OpenStudy (anonymous):

HI everyone i'm new here ! please help me ! can anyone solve this lim? lim[2sinx]=

zepdrix (zepdrix):

Hey there! :) \(\Large\bf \color{#CC0033}{\text{Welcome to OpenStudy! :)}}\)

zepdrix (zepdrix):

What is the limit approaching? Infinity? zero?

zepdrix (zepdrix):

And do the brackets mean anything? Greatest integer function? Or you were just using those for normal brackets? :)

OpenStudy (anonymous):

the main question is about Discontinuity and its usual brackets

OpenStudy (anonymous):

sorry my english is not very good

zepdrix (zepdrix):

\[\large\rm \lim_{x\to?}~2\sin x\]

OpenStudy (anonymous):

its about number of discontinuity of function

zepdrix (zepdrix):

Hmm :d

OpenStudy (anonymous):

you know what i mean?

OpenStudy (anonymous):

Discontinuities in the function

zepdrix (zepdrix):

The limit does not converge as \(\large\rm x\to\infty\) or as \(\large\rm x\to-\infty\). Is that what you mean maybe? :) I don't think we have any discontinuities, because sinx is continuous over all the real numbers. So did you mean diverge instead maybe? :o Hmm thinking..

zepdrix (zepdrix):

There is an "Attach File" button at the bottom of this text box. Maybe you can attach a screenshot of the original problem? :) Or it's not in English? :3

OpenStudy (anonymous):

Chart Dayrhchnd bracket function within half a point of discontinuity ?

OpenStudy (anonymous):

translate it

OpenStudy (anonymous):

i think this is only in our country :)))

OpenStudy (anonymous):

its not enlish

zepdrix (zepdrix):

Oh boy XD

zepdrix (zepdrix):

Hmm, I don't know :c sorry. @ganeshie8 maybe someone can translate and figure this out hehe

OpenStudy (anonymous):

Chart bracket function within a semicircle is a point of continuity (sorry this is correst)

OpenStudy (anonymous):

ok thanks alot